# Recent advances in the numerical solution of multi-order fractional differential equations

### v1

@@ 1 Introduction
The idea of extending the definition of derivatives to the case of a non-integer order of differentiation dates back to 1695, to the correspondence between Leibniz and de l’Hôpital [48], when the meaning of the derivative of order one half was discussed. Although its theoretical foundation is centuries old, the concept of fractional derivative has recently evolved from a purely mathematical curiosity into a vital tool for modeling across a variety of scientific fields (see, e.g., [31, 33, 41, 43, 45, 50, 51], to mention a few), due to its ability to encompass memory terms. In fact, unless the ODE case, the differential operator here considered is of a nonlocal nature [20, 48]. For this reason, the numerical solution of FDEs has been the subject of many researches in the last years, starting from the pioneering work of Lubich [40], where fractional versions of linear multistep formulae are considered: we refer, e.g., to [21, 22, 24, 25, 27, 36, 47], to mention a few of them. More recently, the class of Fractional HBVMs (FHBVMs) [8] has been introduced for the efficient numerical solution of initial value problems of fractional differential equations (FDE-IVPs) in the Caputo form, with one fractional derivative. The basic steps leading to the definition of FHBVMs are the following ones:
  • local expansion of the vector field along a suitable Jacobi polynomial basis, naturally induced by the problem at hand;
@@
Nevertheless, multi-order systems of equations, i.e., systems where different fractional orders appear, are of interest in a number of applications, including material science [52, 48]), population dynamics [4], epidemic modelling [18, 37], pharmacokinetics [55], and chaotic dynamics [5, 30, 38, 49]. In this respect, the extension of FHBVMs for such problems poses important challenges: their study is precisely the goal of this paper.
@@
$y$ $=$ $(y_{1}^{\top},\dots,y_{\nu}^{\top})^{\top}\in\mathbb{R}^{m},\qquad m=\sum_{i=1}^{\nu}m_{i}.$
@@
With this premise, the structure of the paper is as sketched below. At first, in Section 2 we explain the possible approaches for extending FHBVMs to handle problem (1). Later on, in Sections 3 and 4, we provide full details for the corresponding discretization procedure, emphasizing the major difference in the choice of the discrete abscissae and quadratures. The practical solution of the generated discrete problems is then detailed in Section 5. As a result, a new version of the code fhbvm2 in [11], named fhbvm2 2, is given in Section 6, handling at the moment the case $\nu=2$. This latter code is later used in the numerical tests reported in Section 7. At last, some conclusions are given in Section 8.
@@ 2 Fractional HBVMs (FHBVMs)
with $y$ defined according to (1), and $I^{{\alpha}_{i}}f_{i}(y(t))$ the Riemann-Liouville integral corresponding to (2).
@@
$\omega_{i}(c)={\alpha}_{i}(1-c)^{{\alpha}_{i}-1},\quad c\in[0,1],\quad\Longrightarrow\quad\int_{0}^{1}\omega_{i}(c)P_{r}^{i}(c)P_{j}^{i}(c)\mathrm{d}c=\delta_{rj},\quad r,j=0,1,\dots.$ (4)
Next, we expand each vector field $f_{i}$ in (1) along the corresponding family of polynomials, thus obtaining the equivalent problem
$y_{i}^{({\alpha}_{i})}(ch)$ $=$ $f_{i}(y(ch))~\equiv~\sum_{j\geq 0}P_{j}^{i}(c)\gamma_{ij}(y),\qquad c\in[0,1],$ (5)
@@
$y_{i}(h)$ $=$ $T_{i\ell}(h)+\frac{h^{{\alpha}_{i}}}{\Gamma({\alpha}_{i}+1)}\gamma_{i0}(y)~=~T_{i\ell}(h)+\frac{h^{{\alpha}_{i}}}{\Gamma({\alpha}_{i}){\alpha}_{i}}\gamma_{i0}(y)$ (8)
$=$ $T_{i\ell}(h)+\frac{1}{\Gamma({\alpha}_{i})}\int_{0}^{h}(h-x)^{{\alpha}_{i}-1}f_{i}(y(x))\mathrm{d}x$
$\equiv$ $T_{i\ell}(h)+I^{{\alpha}_{i}}f_{i}(y(h)),$
@@
with the Fourier coefficients $\gamma_{ij}(\sigma)$ defined according to (6), by formally replacing $y$ with $\sigma$. As is clear, this is equivalent to requiring the residuals be orthogonal to all polynomials of degree s1s-1, w.r.t. the inner product defined by (4): [^5] This feature is common to HBVMs, obtained when ${\alpha}_{i}=1$ [2, 13, 14].
@@
$\sigma_{i}(h)$ $=$ $T_{i\ell}(h)+\frac{h^{{\alpha}_{i}}}{\Gamma({\alpha}_{i}+1)}\gamma_{i0}(\sigma)$
$=$ $T_{i\ell}(h)+\frac{1}{\Gamma({\alpha}_{i})}\int_{0}^{h}(h-x)^{{\alpha}_{i}-1}f_{i}(\sigma(x))\mathrm{d}x$
@@
$\gamma_{ij}(\sigma^{n})$ $=$ $\int_{0}^{1}\omega_{i}(c)P_{j}^{i}(c)f_{i}(\sigma^{n}(ch_{n}))\mathrm{d}c,\qquad j=0,\dots,s-1,$ (13)
@@
$\sigma^{n}(ch_{n})$ $=$ $\sigma(t_{n-1}+ch_{n}),\qquad c\in[0,1],\qquad n=1,\dots,N.$
@@
For the efficient numerical evaluation of the integrals $I^{{\alpha}_{i}}P_{j}^{i}(c)$ in (7), and $J_{j}^{i}(x)$ in (15), we refer to [9, Section 3.2].
@@
${\bm{1}}_{r}=\left(1,\,\dots,\,1\right)^{\top}\in\mathbb{R}^{r},\qquad{\bm{\alpha}}=({\alpha}_{1},\dots,{\alpha}_{\nu})^{\top}\in\mathbb{R}^{\nu},$
introducing the vectors and functions, with $\otimes$ denoting the Kronecker product,
@@
However, in order to derive a practical numerical method, the $s\nu$ Fourier coefficients (49) (i.e., (50)) need to be approximated by suitable quadrature rules. In the case where in (1) $\nu=1$, this can be efficiently done by using a Gauss-Jacobi quadrature, defined at the zeros of the corresponding $k$th Jacobi polynomial: this results into a FHBVM(k,s)(k,s) method. In the following sections, we study the extension to the case $\nu>1$.
@@ 4 Multiple Orthogonal Polynomials (MOPs)
$({\pi}_{j}(c),{\pi}_{\rho}(c))_{i}$ $=$ $0,\qquad\rho=0,\dots,r,\qquad\qquad i=1,\dots,\mu,$ (73)
$({\pi}_{j}(c),{\pi}_{\rho}(c))_{i}$ $=$ $0,\qquad\rho=0,\dots,r-1,\,\qquad i=\mu+1,\dots,\nu.$
@@
$({\pi}_{i-1}(c),{\pi}_{\rho}(c))_{\iota}$ $=$ $0,\qquad\rho=0,\dots,r-1,\qquad\iota=1,\dots,\eta,$ (78)
@@
$({\pi}_{j-1}(c),{\pi}_{\rho}(c))_{\iota}$ $=$ $0,\qquad\rho=0,\dots,r-1,\qquad\iota=1,\dots,\nu,\qquad\qquad(\mbox{i.e.,}~\eta=\nu).$
@@ 5.1 The case $\nu=1$
In the case $\nu=1$, i.e., when only one fractional derivative occurs, a very efficient Newton-type iteration can be considered, i.e., the so-called blended iteration, at first derived in the ODE case [7, 16, 17, 15, 13], and then adapted for the FDE case [9]. We briefly sketch it, for sake of completeness. In such a case, from (29)–(43), (90)–(98), and (104), with $\nu=1$, considering that $m_{1}=m$, it follows that (106) becomes
@@
$\mbox{solve:}~\left[I_{sm}-h_{n}^{{\alpha}_{1}}X_{11}\otimes f_{1}^{\prime}(\phi^{n}(0))\right]\delta^{j}$ $=$ $-G(\gamma^{j}),$ (109)
$\mbox{update:}~\gamma^{j+1}$ $=$ $\gamma^{j}+\delta^{j},\qquad j=0,1,\dots,$
@@
$\left[I_{sm}-h_{n}^{{\alpha}_{1}}X_{11}\otimes f_{1}^{\prime}(\phi^{n}(0))\right]\in\mathbb{R}^{sm\times sm}.$
@@
$\hat{\eta}_{j}$ $=$ $\rho_{11}(X_{11}^{-1}\otimes I_{m})\eta_{j}$ (121)
$\gamma^{j+1}$ $=$ $\gamma^{j}-\theta^{n}\left[\hat{\eta}_{j}+\theta^{n}\left(\eta_{j}-\hat{\eta}_{j}\right)\right],\qquad j=0,1,\dots.$
@@
$I_{m}-\rho_{11}h_{n}^{{\alpha}_{1}}f_{1}^{\prime}(\phi^{n}(0))\in\mathbb{R}^{m\times m}.$
@@ 5.2 The case $\nu>1$
$\mbox{solve:}~\left[I_{sm}-{\cal P}_{s}^{\top}\Omega\cdot(\,I_{k}\otimes f^{\prime}(\phi^{n}(0))\,)\cdot{\bm{h}}_{n}^{\bm{\alpha}}{\cal I}_{s}\right]{\bm{\delta}}^{j}$ $=$ $-G({\bm{\gamma}}^{j}),$ (122)
$\mbox{update:}~{\bm{\gamma}}^{j+1}$ $=$ ${\bm{\gamma}}^{j}+{\bm{\delta}}^{j},\qquad j=0,1,\dots,$
@@
${\cal P}_{s}^{\top}\Omega\cdot(\,I_{k}\otimes f^{\prime}(\phi^{n}(0))\,)\cdot{\bm{h}}_{n}^{\bm{\alpha}}{\cal I}_{s}\equiv\left(\begin{array}[]{ccc}h_{n}^{{\alpha}_{1}}X_{11}\otimes f_{11}^{n}&\dots&h_{n}^{{\alpha}_{\nu}}X_{1\nu}\otimes f_{1\nu}^{n}\\ \vdots&&\vdots\\ h_{n}^{{\alpha}_{1}}X_{\nu 1}\otimes f_{\nu 1}^{n}&\dots&h_{n}^{{\alpha}_{\nu}}X_{\nu\nu}\otimes f_{\nu\nu}^{n}\end{array}\right)\in\mathbb{R}^{sm\times sm},$
@@
$X_{ij}$ $=$ $\left(\begin{array}[]{ccc}P_{0}^{i}(c_{1})&\dots&P_{0}^{i}(c_{k})\\ \vdots&&\vdots\\ P_{s-1}^{i}(c_{1})&\dots&P_{s-1}^{i}(c_{k})\end{array}\right)\left(\begin{array}[]{ccc}b_{1}^{i}\\ &\ddots\\ &&b_{k}^{i}\end{array}\right)\left(\begin{array}[]{ccc}I^{{\alpha}_{j}}P_{0}^{j}(c_{1})&\dots&I^{{\alpha}_{j}}P_{s-1}^{j}(c_{1})\\ \vdots&&\vdots\\ I^{{\alpha}_{j}}P_{0}^{j}(c_{k})&\dots&I^{{\alpha}_{j}}P_{s-1}^{j}(c_{k})\end{array}\right)$
@@
6 The code fhbvm2 2
[…]

### v2

# A Multi-Order Extension of Fractional HBVMs (FHBVMs)

@@ 1 Introduction
The idea of extending the definition of derivatives to the case of a non-integer order of differentiation dates back to 1695, to the correspondence between Leibniz and de l’Hôpital [57], when the meaning of the derivative of order one half was discussed. Although its theoretical foundation is centuries old, the concept of fractional derivative has recently evolved from a purely mathematical curiosity into a vital tool for modeling across a variety of scientific fields (see, e.g., [38, 40, 48, 52, 54, 59, 60], to mention a few), due to its ability to encompass memory terms. In fact, unless the ODE case, the differential operator here considered is of a nonlocal nature [25, 57]. For this reason, the numerical solution of FDEs has been the subject of many researches in the last years, starting from the pioneering work of Lubich [47], where fractional versions of linear multistep formulae are considered: we refer, e.g., to [26, 27, 31, 32, 34, 43, 56], to mention a few of them. More recently, the class of Fractional HBVMs (FHBVMs) [12] has been introduced for the efficient numerical solution of initial value problems of fractional differential equations (FDE-IVPs) in the Caputo form, with one fractional derivative.
Remark 1
The acronym HBVMs stands for “Hamiltonian Boundary Value Methods”, a class of energy-conserving Runge-Kutta methods for the efficient numerical solution of Hamiltonian ODE problems (see, e.g., the monograph [17]). As matter of fact, when the fractional order of the equation becomes 1 (i.e., the FDE reduces to an ODE), FHBVMs do coincide with HBVMs.
The basic steps leading to the definition of FHBVMs are the following ones:
  • local expansion of the vector field along a suitable Jacobi polynomial basis, naturally induced by the fractional problem at hand (when the equation is an ordinary one, the Jacobi basis reduces to the Legendre polynomial basis, thus resulting into HBVMs);
@@
Nevertheless, multi-order systems of equations, i.e., systems where different fractional orders appear, are of interest in a number of applications, including material science [61, 57]), population dynamics [6], epidemic modelling [20, 44], pharmacokinetics [64], and chaotic dynamics [8, 37, 45, 58]. In this respect, the extension of FHBVMs for such problems poses important challenges. To mention a few of them: the handling of different Jacobi quadratures; a more involved structure and solution of the generated discrete problems; a more involved theoretical analysis of the methods. Their study is precisely the goal of this paper.
@@
$y$ $=$ $(y_{1}^{\top},\dots,y_{\nu}^{\top})^{\top}\in\mathbb{R}^{m},\qquad m=\sum_{i=1}^{\nu}m_{i}.$ (1)
@@
While the numerical solution of linear or nonlinear multi-term FDEs  [^5] I.e., FDEs with different fractional derivatives appearing in the equation. involving Caputo-type fractional derivatives has, over the years, attracted considerable attention by a large number of authors, the general multi-order problem in the form (1), has received significantly less study. Nevertheless, we can point the reader interested in the numerical analysis of such problems to the Adomian decomposition and the variational iteration method studied in [51], the differential transform technique used in [29], the Chebyshev matrix approach implemented in [7, 23], the application of the fractional Jacobi-Picard iteration method in [5], wavelet-type solutions in [21, 65], and the trapezoidal and product integration methods in [34]. We also note that there exist some prior numerical methods implementing a spectral-type approach, with basis functions consisting of fractional-order Jacobi functions [10, 30] or Chelyshkov polynomials [2].
With this premise, the structure of the paper is as follows. At first, in Section 2 we explain the possible approaches for extending FHBVMs to handle problem (1). Later on, in Sections 3 and 4, we provide full details for the corresponding discretization procedure, emphasizing the major difference in the choice of the discrete abscissae and quadratures. The practical solution of the generated discrete problems is then detailed in Section 5, and the stability and accuracy analysis of the methods is given in Section 6. A new version of the code fhbvm2 in [15], named fhbvm2 2, is then presented in Section 7, handling at the moment the case $\nu=2$. This latter code is later used in the numerical tests reported in Section 8. At last, some conclusions are given in Section 9.
@@ 2 Fractional HBVMs (FHBVMs)
with $y$ defined according to (1), and
$I^{{\alpha}_{i}}f_{i}(y(t))=\frac{1}{\Gamma({\alpha}_{i})}\int_{0}^{t}(t-x)^{{\alpha}_{i}-1}f(_{i}(y(x))\mathrm{d}x$
the Riemann-Liouville integral corresponding to (2).
@@
$\omega_{i}(c)={\alpha}_{i}(1-c)^{{\alpha}_{i}-1},\quad c\in[0,1],\quad\Longrightarrow\quad\int_{0}^{1}\omega_{i}(c)P_{r}^{i}(c)P_{j}^{i}(c)\mathrm{d}c=\delta_{rj},\quad r,j=0,1,\dots,$ (4)
with, as is usual, $\delta_{rj}$ denoting the Kronecker symbol. Next, we expand each vector field $f_{i}$ in (1) along the corresponding family of polynomials, thus obtaining the equivalent problem
$y_{i}^{({\alpha}_{i})}(ch)$ $=$ $f_{i}(y(ch))~\equiv~\sum_{j\geq 0}P_{j}^{i}(c)\gamma_{ij}(y),\qquad c\in[0,1],$
@@
$y_{i}(h)$ $=$ $T_{i\ell}(h)+\frac{h^{{\alpha}_{i}}}{\Gamma({\alpha}_{i}+1)}\gamma_{i0}(y)~=~T_{i\ell}(h)+\frac{h^{{\alpha}_{i}}}{\Gamma({\alpha}_{i}){\alpha}_{i}}\gamma_{i0}(y)$
$=$ $T_{i\ell}(h)+\frac{1}{\Gamma({\alpha}_{i})}\int_{0}^{h}(h-x)^{{\alpha}_{i}-1}f_{i}(y(x))\mathrm{d}x~\equiv~T_{i\ell}(h)+I^{{\alpha}_{i}}f_{i}(y(h)),$
@@
with the Fourier coefficients $\gamma_{ij}(\sigma)$ defined according to (6), by formally replacing $y$ with $\sigma$. As is clear, this is equivalent to requiring the residuals be orthogonal to all polynomials of degree s1s-1, w.r.t. the inner product defined by (4): [^6] This feature is common to HBVMs, obtained when ${\alpha}_{i}=1$ [4, 17].
@@
$\sigma_{i}(h)$ $=$ $T_{i\ell}(h)+\frac{h^{{\alpha}_{i}}}{\Gamma({\alpha}_{i}+1)}\gamma_{i0}(\sigma)~=~T_{i\ell}(h)+\frac{1}{\Gamma({\alpha}_{i})}\int_{0}^{h}(h-x)^{{\alpha}_{i}-1}f_{i}(\sigma(x))\mathrm{d}x$
@@
$\gamma_{ij}(\sigma^{n})$ $=$ $\int_{0}^{1}\omega_{i}(c)P_{j}^{i}(c)f_{i}(\sigma^{n}(ch_{n}))\mathrm{d}c,\qquad j=0,\dots,s-1,$
@@
$\sigma^{n}(ch_{n})$ $=$ $\sigma(t_{n-1}+ch_{n}),\qquad c\in[0,1],\qquad n=1,\dots,N.$ (12)
@@
For the efficient numerical evaluation of the integrals $I^{{\alpha}_{i}}P_{j}^{i}(c)$ in (7), and $J_{j}^{i}(x)$ in (14), we refer to [13, Section 3.2].
@@
${\bm{1}}_{r}=\left(1,\,\dots,\,1\right)^{\top}\in\mathbb{R}^{r},\qquad{\bm{\alpha}}=({\alpha}_{1},\dots,{\alpha}_{\nu}),$
introducing the vectors, with $\otimes$ denoting the Kronecker product,
@@
However, in order to derive a practical numerical method, the $s\nu$ Fourier coefficients (48) (i.e., the block vector (49)) need to be approximated by suitable quadrature rules. In the case where in (1) $\nu=1$, this can be efficiently done by using one Gauss-Jacobi quadrature, defined at the zeros of the corresponding $k$th Jacobi polynomial: this results into a FHBVM(k,s)(k,s) method. In the following sections, we study the extension to the case $\nu>1$.
@@ 4 Multiple Orthogonal Polynomials (MOPs)
$({\pi}_{j}(c),{\pi}_{\rho}(c))_{i}$ $=$ $0,\qquad\rho=0,\dots,r,\qquad\qquad i=1,\dots,\mu,$
$({\pi}_{j}(c),{\pi}_{\rho}(c))_{i}$ $=$ $0,\qquad\rho=0,\dots,r-1,\,\qquad i=\mu+1,\dots,\nu.$ (72)
@@
$({\pi}_{i-1}(c),{\pi}_{\rho}(c))_{\iota}$ $=$ $0,\qquad\rho=0,\dots,r-1,\qquad\iota=1,\dots,\eta,$
@@
$({\pi}_{j-1}(c),{\pi}_{\rho}(c))_{\iota}$ $=$ $0,\qquad\rho=0,\dots,r-1,\qquad\iota=1,\dots,\nu,\qquad\qquad(\mbox{i.e.,}~\eta=\nu).$ (77)
@@ 5.1 The case $\nu=1$
In the case $\nu=1$, i.e., when only one fractional derivative occurs, a very efficient Newton-type iteration can be considered, i.e., the so-called blended iteration, at first derived in the ODE case [19, 18, 17], and then adapted for the FDE case [13]. We briefly sketch it, for sake of completeness. In such a case, from (24)–(46), (89)–(101), and (103), with $\nu=1$, considering that $m_{1}=m$, it follows that (105) becomes
@@
$\mbox{solve:}~\left[I_{sm}-h_{n}^{{\alpha}_{1}}X_{11}\otimes f_{1}^{\prime}(\phi^{n}(0))\right]\delta^{j}$ $=$ $-G(\gamma^{j}),$
$\mbox{update:}~\gamma^{j+1}$ $=$ $\gamma^{j}+\delta^{j},\qquad j=0,1,\dots,$ (108)
@@
$\left[I_{sm}-h_{n}^{{\alpha}_{1}}X_{11}\otimes f_{1}^{\prime}(\phi^{n}(0))\right]~\in~\mathbb{R}^{sm\times sm}.$ (120)
@@
$\hat{\eta}_{j}$ $=$ $\rho_{11}(X_{11}^{-1}\otimes I_{m})\eta_{j}$
$\gamma^{j+1}$ $=$ $\gamma^{j}-\theta^{n}\left[\hat{\eta}_{j}+\theta^{n}\left(\eta_{j}-\hat{\eta}_{j}\right)\right],\qquad j=0,1,\dots.$ (121)
@@
$I_{m}-\rho_{11}h_{n}^{{\alpha}_{1}}f_{1}^{\prime}(\phi^{n}(0))~\in~\mathbb{R}^{m\times m},$
@@ 5.2 The case $\nu>1$
$\mbox{solve:}~\left[I_{sm}-{\cal P}_{s}^{\top}\Omega\cdot(\,I_{k}\otimes f^{\prime}(\phi^{n}(0))\,)\cdot{\bm{h}}_{n}^{\bm{\alpha}}{\cal I}_{s}\right]{\bm{\delta}}^{j}$ $=$ $-G({\bm{\gamma}}^{j}),$
$\mbox{update:}~{\bm{\gamma}}^{j+1}$ $=$ ${\bm{\gamma}}^{j}+{\bm{\delta}}^{j},\qquad j=0,1,\dots,$ (122)
@@
${\cal P}_{s}^{\top}\Omega\cdot(\,I_{k}\otimes f^{\prime}(\phi^{n}(0))\,)\cdot{\bm{h}}_{n}^{\bm{\alpha}}{\cal I}_{s}~\equiv~\left(\begin{array}[]{ccc}h_{n}^{{\alpha}_{1}}X_{11}\otimes f_{11}^{n}&\dots&h_{n}^{{\alpha}_{\nu}}X_{1\nu}\otimes f_{1\nu}^{n}\\ \vdots&&\vdots\\ h_{n}^{{\alpha}_{1}}X_{\nu 1}\otimes f_{\nu 1}^{n}&\dots&h_{n}^{{\alpha}_{\nu}}X_{\nu\nu}\otimes f_{\nu\nu}^{n}\end{array}\right)\in\mathbb{R}^{sm\times sm},$
@@
$X_{ij}$ $=$ $\left(\begin{array}[]{ccc}P_{0}^{i}(c_{1})&\dots&P_{0}^{i}(c_{k})\\ \vdots&&\vdots\\ P_{s-1}^{i}(c_{1})&\dots&P_{s-1}^{i}(c_{k})\end{array}\right)\left(\begin{array}[]{ccc}b_{1}^{i}\\ &\ddots\\ &&b_{k}^{i}\end{array}\right)\left(\begin{array}[]{ccc}I^{{\alpha}_{j}}P_{0}^{j}(c_{1})&\dots&I^{{\alpha}_{j}}P_{s-1}^{j}(c_{1})\\ \vdots&&\vdots\\ I^{{\alpha}_{j}}P_{0}^{j}(c_{k})&\dots&I^{{\alpha}_{j}}P_{s-1}^{j}(c_{k})\end{array}\right)$ (133)
@@
6 Analysis of the methods
In this section we sketch a linear stability analysis of FHBVMs and study their convergence order. We also report a few numerical tests to assess the theoretical findings.
6.1 Linear stability analysis
We consider problem (1) in the linear and autonomous case, assuming, without loosing generality, that $\nu=m$:
$y_{i}^{({\alpha}_{i})}(t)=\sum_{j=1}^{m}a_{ij}y_{j}(t),\qquad t>0,\qquad y_{i}(0)=y_{i0}\in\mathbb{R},\qquad 0<{\alpha}_{i}<1,\qquad i=1,\dots,m.$ (136)
We also consider the corresponding shifted orthonormal Jacobi polynomial bases (4). Then, fixing a suitable $s\geq 1$, according to (8)–(9), a FHBVM, used with timestep $t$, basically determines approximations:
$\sigma_{i}^{({\alpha}_{i})}(ct)=\sum_{\nu=0}^{s-1}P_{\nu}^{i}(c)\gamma_{i\nu},\qquad\sigma_{i}(ct)=y_{i0}+t^{{\alpha}_{i}}\sum_{\nu=0}^{s-1}I^{{\alpha}_{i}}P_{\nu}^{i}(c)\gamma_{i\nu},\qquad c\in[0,1],$ (137)
with the coefficients $\{\gamma_{i\nu}\}$ determined such that
$\int_{0}^{1}\omega_{i}(c)P_{\iota}^{i}(c)\left[\sigma_{i}^{({\alpha}_{i})}(ct)-\sum_{j=1}^{m}a_{ij}\sigma_{j}(ct)\right]\mathrm{d}c=0,\qquad\iota=0,\dots,s-1,\quad i=1,\dots,m.$ (138)
From (4) and (137), one then obtains that, recalling that $\delta_{\iota 0}$ denotes the Kronecker symbol,
$\gamma_{i\iota}-\sum_{j=1}^{m}a_{ij}\left[\delta_{\iota 0}y_{j0}+t^{{\alpha}_{j}}\sum_{\nu=0}^{s-1}\left(\int_{0}^{1}\omega_{i}(c)P_{\iota}^{i}(c)I^{{\alpha}_{j}}P_{\nu}^{j}(c)\mathrm{d}c\right)\gamma_{j\nu}\right]=0,\quad\iota=0,\dots,s-1,\quad i=1,\dots,m.$ (139)
More compactly, by setting
$x_{ij}^{\iota\nu}=\int_{0}^{1}\omega_{i}(c)P_{\iota}^{i}(c)I^{{\alpha}_{j}}P_{\nu}^{j}(c)\mathrm{d}c,\qquad i,j=1,\dots,m,\quad\iota,\nu=0,\dots,s-1,$ (140)
and
$\psi_{i}=\sum_{j=1}^{m}a_{ij}y_{j0},\qquad i=1,\dots,m,$ (141)
one derives that (139) is equivalent to:
$\gamma_{i\iota}-\sum_{j=1}^{m}a_{ij}t^{{\alpha}_{j}}\sum_{\nu=0}^{s-1}x_{ij}^{\iota\nu}\gamma_{\nu}^{j}=\delta_{\iota 0}\psi_{i},\qquad\iota=0,\dots,s-1,\quad i=1,\dots,m.$ (142)
Setting $e_{1}\in\mathbb{R}^{s}$ the first unit vector, the vectors
$\gamma^{i}=\left(\begin{array}[]{c}\gamma_{i0}\\ \vdots\\ \gamma_{i,s-1}\end{array}\right)\in\mathbb{R}^{s},\quad i=1,\dots,m,$ (143)
the matrices
$X_{ij}=\left(\begin{array}[]{ccc}x_{ij}^{0,0}&\dots&x_{ij}^{0,s-1}\\ \vdots&&\vdots\\ x_{ij}^{s-1,0}&\dots&x_{ij}^{s-1,s-1}\end{array}\right),\quad B_{ij}=a_{ij}X_{ij}~\in~\mathbb{R}^{s\times s},\qquad i,j=1,\dots,m,$ (144)
and recalling $I_{s}\in\mathbb{R}^{s\times s}$ the identity matrix, one has that (142) can be cast in vector form as
$\left(\begin{array}[]{cccc}I_{s}-B_{11}t^{{\alpha}_{1}}&-B_{12}t^{{\alpha}_{2}}&\dots&-B_{1m}t^{{\alpha}_{m}}\\ -B_{21}t^{{\alpha}_{1}}&I_{s}-B_{22}t^{{\alpha}_{2}}&\dots&-B_{2m}t^{{\alpha}_{m}}\\ \vdots&&\ddots&\vdots\\ -B_{m1}t^{{\alpha}_{1}}&-B_{m2}t^{{\alpha}_{2}}&\dots&I_{s}-B_{mm}t^{{\alpha}_{m}}\end{array}\right)\left(\begin{array}[]{c}\gamma^{1}\\ \gamma^{2}\\ \vdots\\ \gamma^{m}\end{array}\right)=\left(\begin{array}[]{c}e_{1}\otimes\psi_{1}\\ e_{1}\otimes\psi_{2}\\ \vdots\\ e_{1}\otimes\psi_{m}\end{array}\right).$ (145)
Remark 7
Discretizing the integrals in (140) with a quadrature over $k$ points (e.g., $k$ Gauss-Piñeiro abscissae corresponding to the given $\{{\alpha}_{i}\}$) results into a FHBVM(k,s)(k,s) method. In such a case, the matrices $X_{ij}$ defined in (144) do coincide with the matrices $X_{ij}$ defined in (133). Conversely, if a different Gauss-Jacobi quadrature is used for each equation, then $X_{ij}$ reduce to $X_{ij}^{i}$, as defined in Remark 6.
Considering that the functions $\sigma_{i}(ct)$ in (137) can be rewritten as
$\sigma_{i}(ct)=y_{i0}+\sum_{\nu=0}^{s-1}I^{{\alpha}_{i}}P_{\nu}^{i}(c)\left(t^{{\alpha}_{i}}\gamma_{i\nu}\right),\qquad c\in[0,1],\qquad i=1,\dots,m,$
it makes sense to reformulate problem (145) in the equivalent form:
$M_{s}^{\bm{\alpha}}(t)\left(\begin{array}[]{c}t^{{\alpha}_{1}}\gamma^{1}\\ t^{{\alpha}_{2}}\gamma^{2}\\ \vdots\\ t^{{\alpha}_{m}}\gamma^{m}\end{array}\right)=\left(\begin{array}[]{c}e_{1}\otimes\psi_{1}\\ e_{1}\otimes\psi_{2}\\ \vdots\\ e_{1}\otimes\psi_{m}\end{array}\right),$ (146)
with  ${\bm{\alpha}}=({\alpha}_{1},\dots,{\alpha}_{m})$  and
$M_{s}^{\bm{\alpha}}(t)=\left(\begin{array}[]{cccc}t^{-{\alpha}_{1}}I_{s}-B_{11}&-B_{12}&\dots&-B_{1m}\\ -B_{21}&t^{-{\alpha}_{2}}I_{s}-B_{22}&\dots&-B_{2m}\\ \vdots&&\ddots&\vdots\\ -B_{m1}&-B_{m2}&\dots&t^{-{\alpha}_{m}}I_{s}-B_{mm}\end{array}\right).$ (147)
Consequently,
$\left(\begin{array}[]{c}t^{{\alpha}_{1}}\gamma^{1}\\ t^{{\alpha}_{2}}\gamma^{2}\\ \vdots\\ t^{{\alpha}_{m}}\gamma^{m}\end{array}\right)=M_{s}^{\bm{\alpha}}(t)^{-1}\left(\begin{array}[]{c}e_{1}\otimes\psi_{1}\\ e_{1}\otimes\psi_{2}\\ \vdots\\ e_{1}\otimes\psi_{m}\end{array}\right),$ (148)
Considering that the r.h.s. in (146) does not depend on $t$, from (148) one deduces that a FHBVM(k,s)(k,s) method can be called $A$-stable if, for all matrices  $A=(a_{ij})\in\mathbb{R}^{m\times m}$  and  ${\bm{\alpha}}=({\alpha}_{1},\dots,{\alpha}_{m})$  such that the solution of problem (136) is asymptotically stable, one has that matrix
$M_{s}^{\bm{\alpha}}(t)^{-1}D_{s},\qquad{\rm with}\qquad D_{s}=\mathrm{diag}(u\otimes e_{1}),\qquad u=(1,\dots,1)^{\top}\in\mathbb{R}^{m},$
has a bounded norm for all $t>0$ (the reason for introducing matrix $D_{s}$ is due to the fact that, because of the structure of the r.h.s., only the columns $1,s+1,2s+1,\dots,(m-1)s+1$ of the inverse are actually involved). In this case, in fact, it follows that:
$\exists\,K>0~s.t.~\forall\,i=1,\dots,m,~t>0,\mbox{~and~}c\in[0,1]:~|\sigma_{i}(ct)|\leq K.$
As is clear, in general it is not possible to directly assess the $A$-stability of FHBVMs (as it has been done in the case of only one fractional order in [15]). Instead, we provide an example showing that FHBVM(30,s)(30,s), $s=1,\dots,22$, behaves correctly when solving the linear problem
$y_{1}^{(0.5)}(t)=0.1y_{1}(t)+y_{2}(t),\quad y_{2}^{(0.7)}(t)=-y_{1}(t)+0.1y_{2}(t),\quad t>0,\qquad y_{1}(0)=y_{2}(0)=1,$ (149)
whose solution, according to the criteria given, e.g., in [24] (see also [11]), is asymptotically stable. [^14] In fact, in the multi-order case, the much simpler criterion in [49], valid in the single-order case, cannot be used. In more detail, in Figure 1 we plot the function
$\Sigma_{s}(t):=\|(M_{s}^{(0.5,0.7)}(t))^{-1}D_{s}\|,\qquad t\in[10^{-3},10^{4}],\qquad s=1,\dots,22.$ (150)
Figure 1: Plot of the function $\Sigma_{s}(t)$ defined in (150), $s=1,\dots,22$.
From the plots, one deduces that all methods behave in the correct way.
6.2 Convergence analysis
In the sequel (see also [15]), we shall consider the following meshes: